# ToxVidLLM: A Multimodal LLM-based Framework for Toxicity Detection in Code-Mixed Videos

### v1

@@ Abstract
footnotetext: Denotes an equal contribution to this work by the respective authors and are jointly the first authors.
In an era of rapidly evolving internet technology, the surge in multimodal content, including videos, has expanded the horizons of online communication. However, the detection of toxic content in this diverse landscape, particularly in low-resource code-mixed languages, remains a critical challenge. While substantial research has addressed toxic content detection in textual data, the realm of video content, especially in non-English languages, has been relatively underexplored. This paper addresses this research gap by introducing a benchmark dataset, the first of its kind, consisting of 931 videos with 4021 code-mixed Hindi-English utterances collected from YouTube. Each utterance within this dataset has been meticulously annotated for toxicity, severity, and sentiment labels. We have developed an advanced Multimodal Multitask framework built for Toxicity detection in Video Content by leveraging Large Language Models (LLMs), crafted for the primary objective along with the additional tasks of conducting sentiment and severity analysis. ToxVidLLM incorporates three key modules – the Encoder module, Cross-Modal Synchronization module, and Multitask module – crafting a generic multimodal LLM customized for intricate video classification tasks. Our experiments reveal that incorporating multiple modalities from the videos substantially enhances the performance of toxic content detection by achieving an Accuracy and Weighted F1 score of 94.29% and 94.35%, respectively. The code and dataset will be made available at https://github.com/justaguyalways/ToxVidLLM_ACL_2024
@@ 1 Introduction
In an age where social media platforms empower users to become content creators, the digital landscape has witnessed an unprecedented proliferation of information dissemination. By 2023, it’s estimated that 82% of internet traffic will be video content Wilson (2022). As a result, platforms like YouTube and Dailymotion have become major sources of information. A remarkable statistic underscores the colossal impact of these platforms: on YouTube alone, users collectively view more than a billion hours of video content each day [^2] https://blog.youtube/press/. The viral nature of video content is a double-edged sword: it facilitates rapid news propagation yet simultaneously accelerates the dissemination of toxic speech. We adhere to the definition of toxic speech provided by Dixon et al. (2018), which characterizes it as "discourteous, disrespectful, or unreasonable language likely to compel someone to exit a discussion".
@@
Current research in the domain of toxic speech detection primarily focuses on text-based models Kennedy et al. (2020); Roy and Mali (2022); Obaid et al. (2023); Maity et al. (2022b); Das et al. (2022); Maity et al. (2023), with limited exploration of image-based methodologies Yang et al. (2019); Gomez et al. (2020); Kiela et al. (2020); Maity et al. (2022a) and very few works on video data Wu and Bhandary (2020); Rana and Jha (2022); Das et al. (2023); Jha et al. (2024). Detecting harmful actions in videos requires the fusion of multi-frame video and speech processing signals, making direct adaptation of image-based hate detection methods inadequate. Existing toxic content detection methods predominantly rely on text-based modalities, with limited exploration of video content and a focus on monolingual languages like English. However, the surge in code-mixed language use, especially in multilingual countries like India, where people frequently blend Hindi and English in their communication (known as code-mixing Myers-Scotton (1997)), presents a unique challenge for machine learning tool development, as highlighted by Vyas et al. (2014). While studies have addressed toxic content detection in code-mixed social media texts, a significant research gap remains in code-mixed videos.
@@
ii) We have innovated ToxVidLLM, a multimodal multitask framework for detecting toxic videos and analyzing their sentiment and severity. ToxVidLLM integrates three key modules: the Encoder module, the Cross-Modal Synchronization Module, and the Multitask module, to create a versatile Multimodal LLM tailored for video classification tasks. Our framework incorporates a sophisticated gated modality fusion mechanism, empirically proven to outperform standard fusion techniques in ablation studies. We propose a method for synchronizing the text modality with other modalities, yielding promising results as demonstrated in our studies. Notably, our framework is adaptable to various publicly available pre-trained models, serving as modality encoders, making it applicable to diverse problem statements. Our most effective multitask model achieves notable weighted F1-Scores of 94.35%, 86.84%, and 83.42% for Toxicity detection, Severity levels, and Sentiment identification, respectively.
@@ 2 Related Works
While research on the detection of offensive or toxic videos is scarce, particularly in the context of languages such as Portuguese Alcântara et al. (2020), Thai Maity et al. (2024) and English Wu and Bhandary (2020); Rana and Jha (2022); Das et al. (2023), it is worth noting that the existing work in this area predominantly revolves around monolingual languages. The study Maity et al. (2024) leverages textual data to propose a dual-channel network to classify hate and sentiment in low-resource settings. Notably, the two studies Alcântara et al. (2020); Wu and Bhandary (2020) exclusively considered textual features by extracting video transcripts for classification. In contrast, Rana and Jha (2022)’s research takes both textual and audio features for offensive video detection. However, this study confronts issues related to dataset accessibility, insufficiently detailed data curation and annotation processes, and a lack of precise dataset statistics. Das et al. (2023) developed a more comprehensive approach by integrating all three modalities (text, Image, and audio) for hate video detection in English.
@@ 3.1 Data Annotation
Annotation training: Three PhD scholars oversaw the annotation process, well-versed in toxic and offensive content, and the actual annotations were conducted by three undergraduate students proficient in both Hindi and English. Initially, we hired a group of master’s students in linguistics who volunteered via our department email list and compensated them with gift vouchers and an honorarium. To train our annotators, we required gold standard samples with annotations for toxicity, severity, and sentiment labels. Our expert annotators randomly selected 150 samples (a small video of one sentence) and assigned suitable target classes. We considered two toxicity classes (Non-toxic/toxic), three sentiment classes (positive/neutral/negative), and the severity score on a three-point scale (0, 1, 2) for each video sample. Score 0 signifies that there is no indication of toxicity and 1 indicates that the post contains indications of toxicity. However, they are not severe, and a score of 2 indicates that the post contains strong evidence of toxicity (e.g., physical threats or excitement to commit suicide). Expert annotators engaged in discussions to resolve any differences and created 150 gold-standard samples with rationale and target annotations. These 150 annotated examples were divided into three sets, each containing 50 samples, to facilitate a three-phase training process. After each phase, expert annotators collaborated with novice annotators to correct any inaccuracies in the annotations, and the annotation guidelines were updated as needed. Following the conclusion of the third round of training, the top three annotators were selected to annotate the entire dataset containing 4021 samples.
@@ 4 Methodology
Problem Formulation: We formulate our problem as follows: Given an utterance video clip denoted as V, our task is essentially a classification problem. We aim to determine whether the video contains toxic content, as well as assign sentiment and severity labels to it. Each video, $V$, is expressed as a sequence of frames, $F=\left\{f_{1},f_{2},\ldots f_{n}\right\}$, accompanied by its associated audio $A$, and a transcript of the video, $T=\left\{w_{1},w_{2},\ldots w_{m}\right\}$, which consists of a sequence of words. Our goal is to construct a deep learning-based video classifier, denoted as $C:C(T;F;A)\rightarrow y$, where $y$ signifies the actual label of the video for a given task.
In this section, we comprehensively describe our developed LLM-based multimodal-multitask framework ToxVidLLM (see Figure 2) for toxic video detection and its sentiment and severity analysis. To enhance comprehension of our proposed method, we partition it into three distinct components: namely, the Encoder module, Cross Modal Synchronization Module, and the Multitask module.
Figure 2: Architecture of proposed ToxVidLLM model.
@@ 4.1 Encoder Module
Current large language models (LLMs) exhibit substantial capability but are commonly constrained to processing textual data exclusively. In this section, we elucidate our approach to encoding information from diverse modalities.
@@
Please see Appendix B for more details on audio, video and text encoders used in our study.
4.2 Cross Modal Synchronization Module
Modality encoders are commonly trained in isolation, which may introduce variations in the generated representations. Consequently, it becomes imperative to synchronize these distinct representations within a unified space. In this section, we delineate the methodology adopted for aligning these representations.
Modality Synchronization The synchronization strategy is crafted to effectively link features from multiple modalities (audio/video) with a focus on textual features, given their heightened significance in addressing our specific problem statement. This follows because of the inherent noise in speech and visual signals while compared to text and can be empirically observed in the obtained results. The procedure for modality synchronization is delineated as follows:
(1) Encoding: Initially, we utilized pre-trained models such as VideoMAE and Whisper to encode the input video ($V$) and audio ($A$), respectively, as outlined below:
@@ 4.2 Cross Modal Synchronization Module
(2) Abstract Feature Extraction: To mitigate computational expenses and limit the token count in the prefix, we utilize a 1-D convolutional layer (Conv) to compress the length of multi-modal features to a condensed and consistent value. Following this, a linear layer (FC) is employed to modify the hidden size of the features, aligning it with the dimensions of the token embeddings in the LLMs, as described below:
@@
where $C_{v}\in\mathbb{R}^{SL^{{}^{\prime}}\times d_{t}}$ and $C_{a}\in\mathbb{R}^{SL^{{}^{\prime}}\times d_{t}}$ are the abstract features with a fixed length of $SL^{{}^{\prime}}$. Here, $d_{t}$ represents the dimensionality of the embedding matrix $E_{llm}\in\mathbb{R}^{V\times d_{t}}$ associated with the textual LLMs (i.e., HingRoberta or HingGPT in our work) with vocabulary size V.
@@
In our research, we view the attenuated representations of visual ($C_{v}^{s}$) and audio ($C_{a}^{s}$) modalities derived from Equation 4 as the soft tokens utilized by the LLM, acting as the input to the Multitask module.
@@ 4.3 Multitask Module
Typically, the LLM model processes the input transcript $T$, generating a text token embedding $E_{t}$ with dimensions $SL_{t}\times d_{t}$. Here, $SL_{t}$ is the maximum sequence length of the transcript, and $d_{t}$ is the embedding dimension. Here we have added the joint multimodal soft tokens ($J_{va}$) obtained from the Cross-Modal Synchronization Module, appended with the text tokens separated by the special token [SEP], thereby creating a multimodal input to the LLM for a more comprehensive understanding of the input video. Subsequently, the sequence output from the LLM undergoes averaging and is passed through three task-specific fully connected layers, followed by an output softmax layer, facilitating the concurrent solution of three tasks: toxicity, severity, and sentiment detection from a video.
@@ 5 Experimental Results and Analysis
Table 3: Results of proposed ToxVidLLM framework with different modality configurations for three tasks (Toxicity, Severity and Sentiment classification) in single and multitask settings; TF - Timesformer, VM - VideoMAE, WP - Whisper, M-BERT - HingMBERT, GPT2 - HingGPT, RT - HingRoberta.
@@ 5.1 Findings from Experiments
Table 2 shows the results of toxicity, severity and sentiment classification tasks with different baseline models. Results of the proposed ToxVidLLM model are shown in Table 3. From all these reported results, we can conclude the following:
@@
(4) In both single-task and multitask scenarios, our proposed model (ToxVidLLM) consistently outperforms all baselines by a substantial margin. In the three-task settings, the (RT + VM + WP) variants exhibit enhancements in accuracy, surpassing the best baseline model by 6.21%, 8.46%, and 6.61% for toxicity, severity, and sentiment tasks, respectively. These notable improvements underscore the effectiveness of our proposed model, attributed to the incorporation of an innovative modality synchronization module compared to baseline approaches.
@@ 5.1.1 Ablation Study
Table 4: Ablation study to show the effect of gated fusion (GF), multi-head cross attention (MHCA) in proposed ToxVidLLM model
@@
We conducted an ablation study (see Table 4) on our proposed model, ToxVidLLM, to elucidate the impact of gated fusion (GF) and multi-head cross attention (MHCA) in toxic video detection. The removal of GF from ToxVidLLM results in a discernible decrease of 1.72%, 2.73%, and 3.13% in F1-score for toxicity, severity, and sentiment tasks, respectively. This decline underscores the crucial role of the gated fusion module in effectively fusing video and audio modalities, thereby enhancing overall performance. Upon removing the MHCA component from ToxVidLLM, a substantial performance drop is observed across all tasks, affirming the significant impact of MHCA in our proposed model. This proves MHCA is instrumental in generating text-guided audio and video features. Notably, the simultaneous exclusion of both GF and MHCA components results in a substantial drop of 6.63% in the F1 score for the toxicity task, with similar performance reductions observed in other tasks. This considerable decline underscores the pivotal role of the cross-modal synchronization module, emphasizing its capacity to align representations from three distinct modalities within a unified space. Please see the qualitative analysis of our proposed framework in Appendix A.

### v2

# ToxVidLM: A Multimodal Framework for Toxicity Detection in Code-Mixed Videos

@@ Abstract
footnotetext: Denotes an equal contribution to this work by the respective authors and are jointly the first authors.footnotetext: This work has been accepted as a long paper at ACL 2024 Findings.
In an era of rapidly evolving internet technology, the surge in multimodal content, including videos, has expanded the horizons of online communication. However, the detection of toxic content in this diverse landscape, particularly in low-resource code-mixed languages, remains a critical challenge. While substantial research has addressed toxic content detection in textual data, the realm of video content, especially in non-English languages, has been relatively underexplored. This paper addresses this research gap by introducing a benchmark dataset, the first of its kind, consisting of 931 videos with 4021 code-mixed Hindi-English utterances collected from YouTube. Each utterance within this dataset has been meticulously annotated for toxicity, severity, and sentiment labels. We have developed an advanced Multimodal Multitask framework built for Toxicity detection in Video Content by leveraging Language Models (LMs), crafted for the primary objective along with the additional tasks of conducting sentiment and severity analysis. ToxVidLM incorporates three key modules – the Encoder module, Cross-Modal Synchronization module, and Multitask module – crafting a generic multimodal LM customized for intricate video classification tasks. Our experiments reveal that incorporating multiple modalities from the videos substantially enhances the performance of toxic content detection by achieving an Accuracy and Weighted F1 score of 94.29% and 94.35%, respectively. The code and dataset will be made available at https://github.com/justaguyalways/ToxVidLM_ACL_2024
@@ 1 Introduction
In an age where social media platforms empower users to become content creators, the digital landscape has witnessed an unprecedented proliferation of information dissemination. By 2023, it is estimated that 82% of internet traffic will be video content Wilson (2022). As a result, platforms like YouTube and Dailymotion have become major sources of information. A remarkable statistic underscores the colossal impact of these platforms: on YouTube alone, users collectively view more than a billion hours of video content each day [^2] https://blog.youtube/press/. The viral nature of video content is a double-edged sword: it facilitates rapid news propagation yet simultaneously accelerates the dissemination of toxic speech. We adhere to the definition of toxic speech provided by Dixon et al. (2018), which characterizes it as "discourteous, disrespectful, or unreasonable language likely to compel someone to exit a discussion".
@@
Current research in the domain of toxic speech detection primarily focuses on text-based models Kennedy et al. (2020); Roy and Mali (2022); Obaid et al. (2023); Maity et al. (2022b); Das et al. (2022); Maity et al. (2023), with limited exploration of image-based methodologies Yang et al. (2019); Gomez et al. (2020); Kiela et al. (2020); Maity et al. (2022a) and very few works on video data Wu and Bhandary (2020); Rana and Jha (2022); Das et al. (2023); Jha et al. (2024). Detecting harmful actions in videos requires the fusion of multi-frame video and speech processing signals, making direct adaptation of image-based hate detection methods inadequate. Existing toxic content detection methods predominantly rely on text-based modalities, with limited exploration of video content and a focus on monolingual languages like English. However, the surge in code-mixed language use, especially in multilingual countries like India, where people frequently blend Hindi and English in their communication (known as code-mixing Myers-Scotton (1997)), presents a unique challenge for machine learning tool development, as highlighted by Vyas et al. (2014).
While studies have addressed toxic content detection in code-mixed social media texts, a significant research gap remains in code-mixed videos.
@@
ii) We have innovated ToxVidLM, a multimodal multitask framework for detecting toxic videos and analyzing their sentiment and severity. ToxVidLM integrates three key modules: the Encoder module, the Cross-Modal Synchronization Module, and the Multitask module, to create a versatile Multimodal LM tailored for video classification tasks. Our framework incorporates a sophisticated gated modality fusion mechanism, empirically proven to outperform standard fusion techniques in ablation studies. We propose a method for synchronizing the text modality with other modalities, yielding promising results as demonstrated in our studies. Notably, our framework is adaptable to various publicly available pre-trained models, serving as modality encoders, making it applicable to diverse problem statements. Our most effective multitask model achieves notable weighted F1-Scores of 94.35%, 86.84%, and 83.42% for Toxicity detection, Severity levels, and Sentiment identification, respectively.
@@ 2 Related Works
While research on the detection of offensive or toxic videos is scarce, particularly in the context of languages such as Portuguese Alcântara et al. (2020), Thai Maity et al. (2024) and English Wu and Bhandary (2020); Rana and Jha (2022); Das et al. (2023), it is worth noting that the existing work in this area predominantly revolves around monolingual languages. Maity et al. (2024) leverages textual data to propose a dual-channel network to classify hate and sentiment in low-resource settings. Notably, the two studies Alcântara et al. (2020); Wu and Bhandary (2020) exclusively considered textual features by extracting video transcripts for classification. In contrast, Rana and Jha (2022)’s research takes both textual and audio features for offensive video detection. However, this study confronts issues related to dataset accessibility, insufficiently detailed data curation and annotation processes, and a lack of precise dataset statistics. Das et al. (2023) developed a more comprehensive approach by integrating all three modalities (text, Image, and audio) for hate video detection in English.
@@ 3.1 Data Annotation
Annotation training: Three PhD scholars oversaw the annotation process, well-versed in toxic and offensive content, and the actual annotations were conducted by three undergraduate students proficient in both Hindi and English. Initially, we hired a group of masters students in linguistics who volunteered via our department email list and compensated them with gift vouchers and an honorarium. To train our annotators, we required gold standard samples with annotations for toxicity, severity, and sentiment labels. Our expert annotators randomly selected 150 samples (a small video of one sentence) and assigned suitable target classes. We considered two toxicity classes (Non-toxic/toxic), three sentiment classes (positive/neutral/negative), and the severity score on a three-point scale (0, 1, 2) for each video sample. Score 0 signifies that there is no indication of toxicity and 1 indicates that the post contains indications of toxicity. However, they are not severe, and a score of 2 indicates that the post contains strong evidence of toxicity (e.g., physical threats or excitement to commit suicide). Expert annotators engaged in discussions to resolve any differences and created 150 gold-standard samples with rationale and target annotations. These 150 annotated examples were divided into three sets, each containing 50 samples, to facilitate a three-phase training process. After each phase, expert annotators collaborated with novice annotators to correct any inaccuracies in the annotations, and the annotation guidelines were updated as needed. Following the conclusion of the third round of training, the top three annotators were selected to annotate the entire dataset containing 4021 samples.
@@ 4 Methodology
Problem Formulation: We formulate our problem as follows: Given an utterance video clip denoted as V, our task is essentially a classification problem. We aim to determine whether the video contains toxic content, as well as assign sentiment and severity labels to it. Each video, $V$, is expressed as a sequence of frames, $F=\left\{f_{1},f_{2},\ldots f_{n}\right\}$, accompanied by its associated audio $A$ that is sampled at 16kHz to construct a sequence of features $A=\left\{a_{1},a_{2},\ldots a_{l}\right\}$ and a transcript of the video, $T=\left\{w_{1},w_{2},\ldots w_{m}\right\}$, which consists of a sequence of words. Our goal is to construct a deep learning-based video classifier, denoted as $C:C(T;F;A)\rightarrow y$, where $y$ signifies the actual label of the video for a given task.
In this section, we describe our developed LM-based multimodal-multitask framework ToxVidLM (see Figure 2) for toxic video detection and its sentiment and severity analysis. To enhance comprehension of our proposed method, we partition it into three distinct components: namely, the Encoder module, Cross Modal Synchronization Module, and the Multitask module.
Figure 2: Architecture of proposed ToxVidLM model.
@@ 4.1 Encoder Module
Current transformed-based language models (LMs) exhibit substantial capability but are commonly constrained to processing textual data exclusively. In this section, we elucidate our approach to encoding information from diverse modalities.
@@
As per the results obtained, we utilize pre-trained models such as VideoMAE and Whisper to encode the input video ($V$) and audio ($A$) inputs, respectively, as outlined below:
@@
Also, $Z_{v}\in\mathbb{R}^{SL_{v}\times d_{v}}$ and $Z_{a}\in\mathbb{R}^{SL_{a}\times d_{a}}$ respectively. Here, $SL_{v}$ and $SL_{a}$ represent the sequence lengths of video and audio inputs, respectively.$d_{v}$ and $d_{a}$ represent the embedding dimension of encoded audio and video, respectively. Please see Appendix B for more details on audio, video and text encoders used in our study.
4.2 Cross Modal Synchronization Module
Modality encoders are typically trained independently, resulting in discrepancies among the generated representations. Consequently, it becomes imperative to synchronize these distinct representations within a unified space to enhance the overall coherence and effectiveness of multimodal processing. In this section, we present a detailed methodology for aligning these representations.
Modality Synchronization The synchronization strategy aims to effectively correlate features extracted from multiple modalities, such as audio and video, with a primary focus on textual features. This emphasis is due to the heightened significance of textual information in addressing our specific problem statement. Textual features are preferred over auditory and visual signals due to their comparatively lower susceptibility to noise. This empirical observation is consistently reflected in the obtained results, emphasizing the importance of robust textual representation.
The procedure for modality synchronization is delineated as follows:
(1) Abstract Feature Extraction: To mitigate computational expenses and limit the token count in the prefix, we utilize a 1-D convolutional layer (Conv) to compress the length of multi-modal features to a condensed and consistent value. Following this, a linear layer (FC) is employed to modify the hidden size of the features, aligning it with the dimensions of the token embeddings in the LMs, as described below:
@@ 4.2 Cross Modal Synchronization Module
where $C_{v}\in\mathbb{R}^{SL^{{}^{\prime}}\times d_{t}}$ and $C_{a}\in\mathbb{R}^{SL^{{}^{\prime}}\times d_{t}}$ are the abstract features with a fixed length of $SL^{{}^{\prime}}$. Here, $d_{t}$ represents the dimensionality of the embedding matrix $E_{llm}\in\mathbb{R}^{V\times d_{t}}$ associated with the textual LMs (i.e., HingRoberta or HingGPT in our work) with vocabulary size V.
@@
In our research, we view the attenuated representations of visual ($C_{v}^{s}$) and audio ($C_{a}^{s}$) modalities derived from Equation 4 as the soft tokens utilized by the LM, acting as the input to the Multitask module.
@@ 4.3 Multitask Module
Typically, the LM model processes the input transcript $T$, generating a text token embedding $E_{t}$ with dimensions $SL_{t}\times d_{t}$. Here, $SL_{t}$ is the maximum sequence length of the transcript, and $d_{t}$ is the embedding dimension. Here we have added the joint multimodal soft tokens ($J_{va}$) obtained from the Cross-Modal Synchronization Module, appended with the text tokens separated by the special token [SEP], thereby creating a multimodal input to the LM for a more comprehensive understanding of the input video. Subsequently, the sequence output from the LM undergoes averaging and is passed through three task-specific fully connected layers, followed by an output softmax layer, facilitating the concurrent solution of three tasks: toxicity, severity, and sentiment detection from a video.
@@ 5 Experimental Results and Analysis
Table 3: Results of proposed ToxVidLM framework with different modality configurations for three tasks (Toxicity, Severity and Sentiment classification) in single and multitask settings; TF - Timesformer, VM - VideoMAE, WP - Whisper, M-BERT - HingMBERT, GPT2 - HingGPT, RT - HingRoberta.
@@ 5.1 Findings from Experiments
Table 2 shows the results of toxicity, severity and sentiment classification tasks with different baseline models. Results of the proposed ToxVidLM model are shown in Table 3. From all these reported results, we can conclude the following:
@@
(4) In both single-task and multitask scenarios, our proposed model (ToxVidLM) consistently outperforms all baselines by a substantial margin. In the three-task settings, the (RT + VM + WP) variants exhibit enhancements in accuracy, surpassing the best baseline model by 6.21%, 8.46%, and 6.61% for toxicity, severity, and sentiment tasks, respectively. These notable improvements underscore the effectiveness of our proposed model, attributed to the incorporation of an innovative modality synchronization module compared to baseline approaches.
@@ 5.1.1 Ablation Study
Table 4: Ablation study to show the effect of gated fusion (GF), multi-head cross attention (MHCA) in proposed ToxVidLM model
@@
We conducted an ablation study (see Table 4) on our proposed model, ToxVidLM, to elucidate the impact of gated fusion (GF) and multi-head cross attention (MHCA) in toxic video detection. The removal of GF from ToxVidLM results in a discernible decrease of 1.72%, 2.73%, and 3.13% in F1-score for toxicity, severity, and sentiment tasks, respectively. This decline underscores the crucial role of the gated fusion module in effectively fusing video and audio modalities, thereby enhancing overall performance. Upon removing the MHCA component from ToxVidLM, a substantial performance drop is observed across all tasks, affirming the significant impact of MHCA in our proposed model. This proves MHCA is instrumental in generating text-guided audio and video features. Notably, the simultaneous exclusion of both GF and MHCA components results in a substantial drop of 6.63% in the F1 score for the toxicity task, with similar performance reductions observed in other tasks. This considerable decline underscores the pivotal role of the cross-modal synchronization module, emphasizing its capacity to align representations from three distinct modalities within a unified space. Please see the qualitative analysis of our proposed framework in Appendix A.